what is 6 1/4 + 3 1/2

Answer by edjones(8007)

6 1/4 + 3 1/2
=6 1/4 + 3 2/4
4*6+1=25
3*4+2=14
25/4 + 14/4 = 39/4
=9 3/4
.
Ed